Calculate the pressure in atm of 2.50 liters of CH4 gas at STP, if the volume is reduced to 2.00 liters and the temperature is raised to 227oC.
Answer #1

Given:
Pi = 1 atm
Vi = 2.50 L
Vf = 2.00 L
Ti = 273.0 K
Tf = 227.0 oC
= (227.0+273) K
= 500 K

use:
(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)
(1 atm*2.5 L)/(273.0 K) = (Pf*2 L)/(500.0 K)
Pf = 2.29 atm
Answer: 2.29 atm